OpenStudy (anonymous):

How do you find the Taylor series of 1/x when it is centered at 2?

OpenStudy (anonymous):

Without verifying \[\lim_{n \rightarrow \infty} R _{n} (x) = 0\]

OpenStudy (anonymous):

But, I thought you have to use f^(n) (0)?

OpenStudy (anonymous):

That's for the Maclaurin series. Taylor series are more general; they're used for constructing series representations about \(x=c\). If \(c=0\), you have a Maclaurin series.

OpenStudy (anonymous):

Since this one is about \(x=2\), you use \(f^{(n)}(2)\).

OpenStudy (anonymous):

Actually, I made a minor mistake with my series. Should be \[f(x)=\frac{1}{x}\\ f'(x)=-\frac{1}{x^2}\\ f''(x)=\frac{1\cdot2}{x^3}\\ f'''(x)=-\frac{1\cdot2\cdot3}{x^4}\\ ~~~~~~\vdots\\ f^{(n)}(x)=(-1)^n\frac{n!}{x^{n+1}}\] Which gives \[T(x)=\sum_{n=0}^\infty \dfrac{(-1)^n\dfrac{n!}{2^{n+1}}}{n!}(x-2)^n\]

OpenStudy (anonymous):

Equivalently \[T(x)=\sum_{n=0}^\infty \dfrac{(-1)^n}{2^{n+1}}(x-2)^n\] Sorry about the confusion!

OpenStudy (anonymous):

So if you were to write out the first few terms of this series, using the taylor series format, would you use f'(2) for example? @SithsAndGiggles

OpenStudy (anonymous):

Sorry for the delay. Yes, you would use the function's derivatives evaluated at \(x=2\). This is actually where the \(2^{n+1}\) comes from, if you examine the derivation of the nth derivative formula.

OpenStudy (anonymous):

Great!!! Thanks!